OpenStudy (anonymous):

Determine if (-2/3,5) is a solution to -3x+2y=-4. Please show all your work.

OpenStudy (anonymous):

points is(x,y) so put -2/3 into x and 5 into y and check if it's true

OpenStudy (anonymous):

-3(-2/3) +2(5) = 2+10 = 12 and its not equal to -4 so its not a solution.